OpenStudy (anonymous):

(2a^x +1/2 b^x)^2 i need help:(

OpenStudy (anonymous):

(2a^x + 1/2b^x) ^ 2 = (2a^x + 1/2b^x)(2a^x + 1/2b^x) (2ax)(2ax)= 4a^2x^2 (2ax)(1/2bx)=abx^2 (1/2bx)(2ax)=abx^2 (1/2bx)(1/2bx)=0.25b^2x^2 4a^{2}x ^{2}+2abx ^{2}+0.25b ^{2}x ^{2}

OpenStudy (anonymous):

ya a^x

OpenStudy (anonymous):

\[4a ^{2x} + 2 a^x b^x + \frac{ 1 }{ 4 }b ^{2x}\]

OpenStudy (anonymous):

\[2a ^{x} \times 2a ^{x} = 4a ^{2x}\]

OpenStudy (anonymous):

Do you see how to do it now?

OpenStudy (anonymous):

ya thank u so much!